# Accelerating Diffusion Transformers with Dual Feature Caching

### v1

@@ 4.3 Ablation Studies
As shown in Table 5, we compare the Conservative-only, Aggressive-only, and alternating DuCa structures on PixArt-$\alpha$ in terms of FLOPs, FID-30k, and CLIP score. The Conservative-only structure appears to have the best generation quality but incurs an excessively high computational cost. In contrast, Aggressive-only reduces computation to nearly half of Conservative-only but suffers significantly in generation quality. Our proposed DuCa structure achieves a computational load between the two, with only a slight 0.13 increase in FID-30k compared to the high-computation Conservative-only approach, while improving text-image alignment. This indicates that DuCa effectively eliminates most redundant computation present in Conservative-only.
5 Conclusion
@@
This paper first gives an in-depth study of the caching error on aggressive and conservative methods, and then introduces dual caching DuCa, which performs aggressive and conservative alternatively to leverage their advantages in different caching steps. Besides, V-Caching is introduced to perform token-wise conservative caching based on the norm of the value matrix of the tokens, making it compatible with default optimized attention computation such as FlashAttention, allowing feature caching to be practical in real applications. We hope this paper can attract more attention to feature caching for its real benefits, to make it simple but effective, instead of being complex.
Supplementary Material
6 Experimental Details
In this section, more details of the experiments are provided.
6.1 Model Configuration
As mentioned in 4.1.1 , experiments on three models from different tasks, PixArt-$\alpha$[7] for text-to-image generation, OpenSora[4] for text-to-video generation, and DiT[26] for class-conditional image generation, are presented. In this section, a more detailed hyperparameter configuration scheme is provided.
  • PixArt-$\alpha$: The average cache cycle length is set to $\mathcal{N}=3$ with an average caching ratio of $R=70\%$ for ToCa[35], and $\mathcal{N}=3$ with $\mathcal{R}=25\%$ for the DuCa method to maintain a similar acceleration ratio. Additionally, in the FORA [28] scheme, $\mathcal{N}=2$. Conservative caching steps are configured to occur on odd-numbered steps following fresh steps, while aggressive caching steps occur on even-numbered steps following fresh steps. The proportion of conservative caching steps across layers and steps is identical to that in ToCa[35].
  • OpenSora: The average cache cycle length is $\mathcal{N}=3$ for temporal attention, spatial attention, and MLP, and $\mathcal{N}=6$ for cross-attention, with an average caching ratio $\mathcal{R}=85\%$ exclusively for the MLP in ToCa[35] and DuCa, while all other modules use a caching ratio of $100\%$. In the FORA [28]scheme, all modules are set to $\mathcal{N}=2$.
    For $\textbf{PAB}^{1*}$ [33], temporal attention is set to $\mathcal{N}=2$, spatial attention to $\mathcal{N}=4$, cross-attention to $\mathcal{N}=6$, and the MLP module is fully computed. For $\textbf{PAB}^{2*}$ [33], $\mathcal{N}$ is set to 3, 5, and 7 for temporal, spatial, and cross-attention, respectively. For $\textbf{PAB}^{3*}$ [33], $\mathcal{N}$ is set to 5, 7, and 9 for temporal, spatial, and cross-attention, respectively.
    For DuCa, conservative caching steps are set on odd-numbered steps following fresh steps, while aggressive caching steps are set on even-numbered steps following fresh steps. The proportion of conservative caching steps across layers and steps is identical to that in ToCa[35].
  • DiT: The average cache cycle length is set to $\mathcal{N}=4$ with an average caching ratio of $R=93\%$ for ToCa[35](a), $\mathcal{N}=3$ with $R=93\%$ for ToCa[35](b), and $\mathcal{N}=3$ with $R=95\%$ for both DuCa(a) and (b). Additionally, the average cache cycle length for FORA [28]is set to $\mathcal{N}=3$.
    Furthermore, for DuCa(a), conservative caching steps are set on even-numbered steps following fresh steps, while aggressive caching steps are set on odd-numbered steps. In contrast, for DuCa(b), the arrangement is reversed.
    Since the average cache cycle length $\mathcal{N}=3$ here, DuCa(a), which places aggressive caching steps immediately after the fresh step, achieves a higher acceleration ratio compared to DuCa(b), which places conservative caching steps in the first position after the fresh step.
6.2 Anonymous Page for video presentation
To further demonstrate the superiority of DuCa in video generation, we have created an anonymous GitHub page. Please visit https://duca2024.github.io/DuCa/ for a more detailed view. Besides, the videos are also available through the Supplementary Material.

### v2

@@ V-C Ablation Studies
As shown in Table VI, we compare the conservative-only, aggressive-only, and alternating DuCa structures, where the aggressive structure is used in activated odd steps and the conservative structure in even steps, on FLUX.1-dev in terms of Image Reward and FLOPs. The conservative-only structure appears to have good generation quality but incurs an excessively high computational cost. In contrast, aggressive-only reduces computation to nearly half of conservative-only but suffers significantly in generation quality. Our proposed DuCa structure achieves a balanced computational load between the two, reducing the computational cost compared to the conservative-only method while maintaining comparable or even better generation quality.
Choice of Cycle Length and Cache Ratio
We further conduct ablation experiments to investigate the influence of cycle length ($N$) and cache ratio ($R$) on performance, as illustrated in Figure 7. When varying $N$ from 3 to 8, we observe that shorter cycles incur higher FLOPs, while excessively long cycles result in noticeable quality degradation. The setting $N=5$ provides a favorable trade-off, preserving generation quality while reducing computation. Similarly, for cache ratio $R$, lowering $R$ below $0.7$ significantly harms quality, whereas $R=0.9$ eliminates redundant computation without compromising fidelity. Therefore, we adopt $N=5$ and $R=0.9$ as default hyperparameters in our main experiments.
VI Conclusion
@@
This paper first gives an in-depth study of the caching error on aggressive and conservative methods and then introduces dual caching DuCa, which performs aggressive and conservative alternatively to leverage their advantages in different caching steps. Moreover, we have discovered the value of random selection and further explored its implications, revealing that diversity plays a more critical role in token selection than previously thought. Random selection not only demonstrates excellent performance but also seamlessly integrates with efficient attention methods like FlashAttention, making feature caching a practical and viable approach in real-world applications. We hope this paper can attract more attention to feature caching for its real benefits, to make it simple but effective, instead of complex.